OpenStudy (anonymous):

For the following system, if you isolated y in the first equation to use the Substitution Method, what expression would you substitute into the second equation? 2x + y = 8 −x − y = −5 8 + 2x 8 − 2x −8 − 2x −8 + 2x

OpenStudy (cwrw238):

2x + y = 8 subtract 2x from both sides y = 8 - 2x - tags what you substitute into second equation

OpenStudy (cwrw238):

* thats what

OpenStudy (anonymous):

What about for this question? Solve the following system. y = x + 4 5x + y = 16 (2, −6) (−2, 6) (6, 2) (2, 6) would i be doing the same thing? I got (2,-6) idk if thats right @cwrw238

OpenStudy (cwrw238):

well plug in x+4 for y in second equation 5x + x + 4 = 16 6x = 12 x = 2 now sub x= 2 into y= x + 4 gives y = 2+4 = 6 (2,6)

OpenStudy (nali):

@Reemxoxo do u still need help?